40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9640     Entry time: Fri Feb 14 21:03:13 2014
Author: Koji 
Type: Update 
Category: General 
Subject: Y end "BS" 

As I didn't have the green laser PZT feedback for the laser temp control, I went to the yend to check out what's the situation.

I found horrible and disgusting "remnants".

WHAT ARE THESE BSs AT THE Y END?

- The table enclosure was left open

- A (hacky) DB25 cable with clips was blocking the corridor and I was about to trip with the cable.

- This DB25 cable went to the table without going through the air tight feedthrough that is designed for this purpose.

- An SR560 (presumably for the openloop TF measurement) was left inserted in the loop with entangled cables connected to the servo box.

- Of course the laser PZT out mon was left unplugged.

Even after cleaning these cables (a bit), the end setups (including the X end too) are too amature.
Everything is so hacky. We should not allow ourselves to construct this level of setup everytime
we work on any system. This just adds more and more mysteries and eventually we can't handle
the complexity.

ELOG V3.1.3-